# Magnon Blockade with Skyrmion Qubit-Magnon Coupling in a Hybrid Quantum System

### v1

@@ III Magnon Blockade and the steady-state equal-time second-order correlation function
$|\psi\rangle=C_{g0}|g0\rangle+C_{g1}|g1\rangle+C_{e0}|e0\rangle+C_{e1}|e1\rangle+C_{g2}|g2\rangle$
where $C_{ms}$ ($ms=g0,g1,e0,e1,g2$) are the probability amplitudes for the states $|ms\rangle$, and $|g(e),n\rangle$ represents the ground (excited) state of the Skyrmion qubit and the Fock state of the magnon mode, respectively.
@@
$i\dot{C}_{g,0}$ $=\frac{\Omega_{s}}{2}C_{e,0}+\Omega_{d}C_{g,1}$ (5a)
$\vspace{-1em}i\dot{C}_{e,0}$ $=(\Delta_{s}-i\frac{k_{s}}{2})C_{e,0}+\frac{g_{ms}}{2}C_{g,1}+\frac{\Omega_{s}}{2}C_{g,0}$ (5b)
$\vspace{-1em}i\dot{C}_{g,1}$ $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g,1}+\frac{g_{ms}}{2}C_{e,0}+\Omega_{d}C_{g,0}$ (5c)
$\vspace{-1em}i\dot{C}_{e,1}$ $=(\Delta_{m}+\Delta_{s}-i\frac{k_{m}}{2}-i\frac{k_{s}}{2})C_{e,1}+\frac{\sqrt{2}}{2}g_{ms}C_{g,2}$
$\quad+\frac{\Omega_{s}}{2}C_{g,1}+\Omega_{d}C_{e,0}$ (5d)
$\vspace{-1em}i\dot{C}_{g,2}$ $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g,2}+\sqrt{2}\Omega_{d}C_{g,1}+\frac{\sqrt{2}}{2}g_{ms}C_{e,1}\vspace{-1em}$ (5e)
@@
00 0 $=\frac{\Omega_{s}}{2}C_{e,0}+\Omega_{d}C_{g,1},$ (6a)
00 0 $=(\Delta_{s}-i\frac{k_{s}}{2})C_{e,0}+\frac{g_{ms}}{2}C_{g,1}+\frac{\Omega_{s}}{2}C_{g,0},$ (6b)
00 0 $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g,1}+\frac{g_{ms}}{2}C_{e,0}+\Omega_{d}C_{g,0},$ (6c)
00 0 $=(\Delta_{m}+\Delta_{s}-i\frac{k_{m}}{2}-i\frac{k_{s}}{2})C_{e,1}+\frac{\sqrt{2}}{2}g_{ms}C_{g,2}$
$\quad+\frac{\Omega_{s}}{2}C_{g,1}+\Omega_{d}C_{e,0},$ (6d)
00 0 $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g,2}+\sqrt{2}\Omega_{d}C_{g,1}+\frac{\sqrt{2}}{2}g_{ms}C_{e,1}.$ (6e)
In the weak-driving regime, where $\Omega_{d},\Omega_{s}\ll k_{m},k_{s}$, the following relationships hold: $|C_{g,0}|\gg|C_{g,1}|,|C_{g,2}|,\,|C_{e,0}|\gg|C_{g,3}|,|C_{e,1}|,|C_{e,2}|$. By applying perturbation theory, we neglect the higher-order terms that contribute minimally in the equation of motion for the lower-order variables. We can derive the steady-state solutions for the probability amplitudes
$C_{g,0}$ $\approx 1$ (7a)
$C_{e,0}$ $=\frac{-2\Omega_{d}g_{ms}+2\tilde{\Delta}_{m}\Omega_{s}}{-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2}}$ (7b)
$C_{g,1}$ $=\frac{4\tilde{\Delta}_{s}\Omega_{d}-\Omega_{s}g_{ms}}{-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2}}$ (7c)
$C_{e,1}$ $=\frac{-\Omega_{s}DA-2A\Omega_{d}E+8\tilde{\Delta}_{s}\Omega_{d}g_{ms}-2g_{ms}^{2}\Omega_{d}\Omega_{s}}{(-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2})(2AB-g_{ms}^{2})}$ (7d)
$C_{g,2}$ $=\frac{-2\sqrt{2}BC+\Omega_{s}Dg_{ms}+2\Omega_{d}Eg_{ms}}{(-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2})(2\sqrt{2}AB-\sqrt{2}g_{ms}^{2})}$ (7e)
where $\tilde{\Delta}_{m}=\Delta-i\frac{k_{m}}{2},\tilde{\Delta}_{s}=\Delta-i\frac{k_{s}}{2},A=2\left(\Delta-i\frac{k_{m}}{2}\right),B=\left(2\Delta-i\frac{k_{m}}{2}-i\frac{k_{s}}{2}\right),C=4\sqrt{2}\tilde{\Delta}_{s}\Omega_{d}^{2}-\sqrt{2}g_{ms}\Omega_{d}\Omega_{s},D=4\tilde{\Delta}_{s}\Omega_{d}-\Omega_{s}g_{ms},E=-2\Omega_{d}g_{ms}+2\tilde{\Delta}_{m}\Omega_{s}$.
@@
$g^{(2)}(0)$ $\approx\frac{2\left|C_{g2}\right|^{2}}{\left|C_{g1}\right|^{4}}=\frac{2\left|\eta_{1}\right|^{2}\cdot\left|\chi\right|^{2}}{\left|\eta_{2}\right|^{2}\left|D\right|^{4}}$ (8)
@@
III.2 Numerical Simulation of the Second-Order Correlation Function
@@
here, $\kappa_{s}$ and $\kappa_{m}$ represent the decay rates of the skyrmion qubit and the magnon mode, respectively. The Pauli operators are defined as $\tilde{\sigma}_{+}=|\psi_{2}\rangle\langle\psi_{1}|$ and $\tilde{\sigma}_{-}=|\psi_{1}\rangle\langle\psi_{2}|$, where $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$ are the eigenstates of the skyrmion qubit. For simplicity, we assume $\kappa_{s}=\kappa_{m}$, where $\rho$ is the system density matrix. The Lindblad superoperators are defined as follows
@@
$g^{(2)}(0)=\frac{\text{Tr}(\rho m^{\dagger}m^{\dagger}mm)}{[\text{Tr}(\rho m^{\dagger}m)]^{2}}=\frac{\langle m^{\dagger}m^{\dagger}mm\rangle}{\langle m^{\dagger}m\rangle^{2}}.$ (12)
The second-order correlation function $g^{(2)}(0)$ can be numerically obtained by solving the master equation given in Eq. (9) using Monte Carlo wave function simulationJohansson et al. (2013), alongside the effective Hamiltonian $H_{\text{eff}}$ from Eq. (2). If $g^{(2)}(0)>1$, this indicates that the steady-state magnon number distribution follows super-Poissonian statistics, exhibiting magnon bunching. Conversely, if $10<g^{(2)}(0)<$ 1, the magnon number distribution satisfies sub-Poissonian statistics, known as magnon antibunching Rabl (2011); Liu et al. (2010); Zheng et al. (2019). Specifically, when $g^{(2)}(0)\rightarrow 0$, magnon blockade occurs. In our scheme, the second-order correlation function can be as low as $g^{(2)}(0)\rightarrow 10^{-6}$, which is a strong indication of magnon blockade Tabuchi et al. (2015); Fujita et al. (2016); Liu et al. (2019). Therefore, our study provides theoretical support for the generation of a single-photon source, which is significant for investigating fundamental aspects of quantum magnons and for potential applications in magnon-based quantum simulation and hybrid quantum devices Li et al. (2019); Hou and Liu (2019); Pan et al. (2024a).
In our simulation, we examine how the equal-time second-order correlation function depends on various system parameters, such as the driving detuning, the effective qubit-magnon coupling strength, the dissipation rates of the magnon mode and the skyrmion-qubit, and the thermal magnon occupation number, decay rate.
@@
Figure 2: The variation of the steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ is plotted varies with (a) as a function of the driving detuning $\Delta/\gamma$ and the skyrmion qubit-magnon coupling strength $g_{ms}/\gamma$, and (b) as a function of the control field strength $\Omega_{s}/\gamma$ and the probe field strength $\Omega_{d}/\gamma$. To maintain consistency throughout the paper, all frequencies are scaled using $\gamma=2\pi\times 1$ MHz. In (a), $\Omega_{s}/\gamma=0.06$ and $\Omega_{d}/\gamma=0.01$, while in (b), $\Delta/\gamma=10$ and $g_{ms}/\gamma=19.6$. The other system parameters are $\kappa_{m}/\gamma=\kappa_{q}/\gamma=1$, and $n_{th}=0$.
@@
In Fig. 2(a), the steady-state second-order correlation function $\log_{10}g^{(2)}(0)$ varies with the drive detuning $\Delta/\gamma$ and the coupling strength $g_{ms}/\gamma$ between the magnon and the skyrmion-qubit. The color gradient from deep purple to yellow represents the transition of $g^{(2)}(0)$ from strong quantum correlations to classical correlations. Near the white dashed lines, the drive detuning is approximately half of the coupling strength, i.e., $\Delta/\gamma\approx\pm\frac{1}{2}g_{ms}/\gamma$, at which point $g^{(2)}(0)$ reaches its minimum, indicating significant antibunching and a pronounced magnon blockade effect in the system. The relationship between the second-order correlation function $g^{(2)}(0)$ and the control field strength $\Omega_{s}/\gamma$, as well as the probe field strength $\Omega_{d}/\gamma$, is shown in Fig. 2(b). As $\Omega_{s}/\gamma$ and $\Omega_{d}/\gamma$ increase, $g^{(2)}(0)$ exhibits significant changes. Notably, $g^{(2)}(0)$ reaches its minimum near the white dashed lines, suggesting the possibility of achieving magnon blockade by appropriately tuning the external driving fields.
@@
Figure 3: The steady-state $\log_{10}g^{(2)}(0)$ is plotted as a function of (a) the coupling strength $g_{ms}$ and (b) the control field strength $\Omega_{s}/\gamma$, as well as the dissipation rate $\kappa$, under the condition $\Delta/\gamma\approx\frac{1}{2}g_{ms}/\gamma$. In (a), $\Omega_{s}/\gamma=0.06$ and $\Omega_{d}/\gamma=0.01$, while in (b) $g_{ms}/\gamma=19.6$, and $n_{th}=0$.
It is essential to clearly state that the detection field and the driving field are indispensable for enhancing magnon blockade. In Fig. 3, we explore the optimal connection between the coupling strength $g_{ms}$, the optimal driving field strength $\Omega_{s}$, and the dissipation rate $\kappa$. In Fig. 3(a), at high coupling strength $g_{ms}$ and low dissipation rate $\kappa$, log10 $g^{(2)}(0)$ remains at a low value (deep purple), with a magnitude less than $10^{-8}$, indicating a strong antibunching effect. We fixed a strong coupling strength $g_{ms}/\gamma=19.6$ and optimized the $g^{(2)}(0)$ function by adjusting the parameter space of $\Omega_{s}$ and $\kappa$, as shown in Fig. 3(b). Our goal is to identify the optimal driving strength and dissipation range under this coupling strength. When $\Omega_{s}/\gamma$ approaches 0.06, log10 $g^{(2)}(0)$ shows a clear low-value region (deep purple band), particularly in the range of $\kappa/\gamma$ from 0 to 0.6, indicating that at this driving field strength, the second-order correlation function reaches its minimum, and the magnon blockade effect is most pronounced. Interestingly, there exists a well-defined optimal driving field strength, detection field strength, and a specific dissipation range to maximize the magnon blockade effect.
@@
Figure 4: The equal-time second-order correlation function $g^{(2)}(0)$ is shown as a function of the normalized detuning $\Delta/\gamma$. As illustrated in (a), for the magnon drive field strengths $\Omega_{d}/\gamma=0.004$, 0.01, and 0.012, the Skyrmion drive field $\Omega_{s}/\gamma$ is kept constant at 0.06. In (b), for the skyrmion drive field strengths $\Omega_{s}/\gamma=0.001$, 0.05, and 0.09, the magnon drive field $\Omega_{d}/\gamma$ is fixed at 0.01. The coupling strength between the magnon and skyrmion modes is $g_{\text{ms}}/\gamma=19.6$. The other system parameters are $\kappa_{m}/\gamma=\kappa_{q}/\gamma=0.15$, and $n_{th}=0$.
Next, we observe that the second-order correlation function $g^{(2)}(0)$ exhibits a strong dependence on the driving detuning, with clearer results presented in Fig. 4. The optimal dissipation region $k$ is selected with $k/\gamma=0.15$. Specifically, we plot the second-order correlation function $g^{(2)}(0)$ as a function of the driving detuning $\Delta/\gamma$. The variation of $g^{(2)}(0)$ with detuning $\Delta/\gamma$ for different magnon driving field strengths ($\Omega_{d}/\gamma=0.004,0.01,0.012$) is shown in Fig. 4(a), where the skyrmion driving field is kept constant at $\Omega_{s}/\gamma=0.06$. As expected, a strong antibunching effect is observed around $\Delta/\gamma\approx 10$, indicating a pronounced magnon blockade effect. In this range, the antibunching effect is sensitive to modulation of the probe and driving fields. However, it is also observed that the antibunching effect around $\Delta/\gamma\approx-10$ is less sensitive to such modulation. When $\Omega_{d}/\gamma=0.01$, the system displays the strongest antibunching effect, with $g^{(2)}(0)$ approaching $10^{-6}$. In Fig. 4(b), the dependence of $g^{(2)}(0)$ on the skyrmion driving field strengths ($\Omega_{s}/\gamma=0.001,0.05,0.09$) is shown, while the magnon driving field remains fixed at $\Omega_{d}/\gamma=0.05$. Similar to the magnon driving fields, the strongest antibunching effect is observed at $\Omega_{s}/\gamma=0.05$, where $g^{(2)}(0)$ approaches $10^{-5}$, indicating optimal suppression of paired magnon excitations.

### v2

@@ III.1 The solution of the second-order correlation function
$|\psi\rangle=$ $\,C_{g^{\prime},0}|g^{\prime},0\rangle+C_{g^{\prime},1}|g^{\prime},1\rangle+C_{e^{\prime},0}|e^{\prime},0\rangle$
$+C_{e^{\prime},1}|e^{\prime},1\rangle+C_{g^{\prime},2}|g^{\prime},2\rangle$ (4)
where $C_{m^{\prime},s}$ ($m=g^{\prime},e^{\prime}\;\text{and}\;s=0,1,2$) are the probability amplitudes for the states $|ms\rangle$, and $|g(e),n\rangle$ represents the ground (excited) state of the skyrmion qubit and the Fock state of the magnon mode, respectively.
@@
$i\dot{C}_{g^{\prime},0}$ $=\frac{\Omega_{s}}{2}C_{e^{\prime},0}+\Omega_{d}C_{g^{\prime},1}$ (5a)
$\vspace{-1em}i\dot{C}_{e^{\prime},0}$ $=(\Delta_{s}-i\frac{k_{s}}{2})C_{e^{\prime},0}+\frac{g_{ms}}{2}C_{g^{\prime},1}+\frac{\Omega_{s}}{2}C_{g^{\prime},0}$ (5b)
$\vspace{-1em}i\dot{C}_{g^{\prime},1}$ $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g^{\prime},1}+\frac{g_{ms}}{2}C_{e^{\prime},0}+\Omega_{d}C_{g^{\prime},0}$ (5c)
$\vspace{-1em}i\dot{C}_{e^{\prime},1}$ $=(\Delta_{m}+\Delta_{s}-i\frac{k_{m}}{2}-i\frac{k_{s}}{2})C_{e^{\prime},1}+\frac{\sqrt{2}}{2}g_{ms}C_{g^{\prime},2}$
$\quad+\frac{\Omega_{s}}{2}C_{g^{\prime},1}+\Omega_{d}C_{e^{\prime},0}$ (5d)
$\vspace{-1em}i\dot{C}_{g^{\prime},2}$ $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g^{\prime},2}+\sqrt{2}\Omega_{d}C_{g^{\prime},1}+\frac{\sqrt{2}}{2}g_{ms}C_{e^{\prime},1}\vspace{-1em}$ (5e)
@@
00 0 $=\frac{\Omega_{s}}{2}C_{e^{\prime},0}+\Omega_{d}C_{g^{\prime},1},$ (6a)
00 0 $=(\Delta_{s}-i\frac{k_{s}}{2})C_{e^{\prime},0}+\frac{g_{ms}}{2}C_{g^{\prime},1}+\frac{\Omega_{s}}{2}C_{g^{\prime},0},$ (6b)
00 0 $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g^{\prime},1}+\frac{g_{ms}}{2}C_{e^{\prime},0}+\Omega_{d}C_{g^{\prime},0},$ (6c)
00 0 $=(\Delta_{m}+\Delta_{s}-i\frac{k_{m}}{2}-i\frac{k_{s}}{2})C_{e^{\prime},1}+\frac{\sqrt{2}}{2}g_{ms}C_{g^{\prime},2}$
$\quad+\frac{\Omega_{s}}{2}C_{g^{\prime},1}+\Omega_{d}C_{e^{\prime},0},$ (6d)
00 0 $=(\Delta_{m}-i\frac{k_{m}}{2})C_{g^{\prime},2}+\sqrt{2}\Omega_{d}C_{g^{\prime},1}+\frac{\sqrt{2}}{2}g_{ms}C_{e^{\prime},1}.$ (6e)
In the weak-driving regime ($\Omega_{d},\Omega_{s}\ll\kappa_{m},\kappa_{s}$), where $|C_{g^{\prime},0}|\gg|C_{g^{\prime},1}|,|C_{g^{\prime},2}|$ and $|C_{e^{\prime},0}|\gg|C_{g^{\prime},3}|,|C_{e^{\prime},1}|,|C_{e^{\prime},2}|$. By applying perturbation theory, we neglect the higher-order terms that contribute minimally in the equation of motion for the lower-order variables. We can derive the steady-state solutions for the probability amplitudes
$C_{g^{\prime},0}$ $\approx 1$ (7a)
$C_{e^{\prime},0}$ $=\frac{-2\Omega_{d}g_{ms}+2\tilde{\Delta}_{m}\Omega_{s}}{-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2}}$ (7b)
$C_{g^{\prime},1}$ $=\frac{4\tilde{\Delta}_{s}\Omega_{d}-\Omega_{s}g_{ms}}{-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2}}$ (7c)
$C_{e^{\prime},1}$ $=\frac{-\Omega_{s}DA-2A\Omega_{d}E+8\tilde{\Delta}_{s}\Omega_{d}g_{ms}-2g_{ms}^{2}\Omega_{d}\Omega_{s}}{(-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2})(2AB-g_{ms}^{2})}$ (7d)
$C_{g^{\prime},2}$ $=\frac{-2\sqrt{2}BC+\Omega_{s}Dg_{ms}+2\Omega_{d}Eg_{ms}}{(-4\tilde{\Delta}_{m}\tilde{\Delta}_{s}+g_{ms}^{2})(2\sqrt{2}AB-\sqrt{2}g_{ms}^{2})}$ (7e)
where $\tilde{\Delta}_{m}=\Delta_{m}-i\frac{k_{m}}{2},\,\tilde{\Delta}_{s}=\Delta_{s}-i\frac{k_{s}}{2},\,A=2\big{(}\Delta_{m}-i\frac{k_{m}}{2}\big{)},\,B=\big{(}\Delta_{m}+\Delta_{s}\big{)}-i\frac{k_{m}}{2}-i\frac{k_{s}}{2},\,C=4\sqrt{2}\tilde{\Delta}_{s}\Omega_{d}^{2}-\sqrt{2}g_{ms}\Omega_{d}\Omega_{s},\,D=4\tilde{\Delta}_{s}\Omega_{d}-\Omega_{s}g_{ms},\,E=-2\Omega_{d}g_{ms}+2\tilde{\Delta}_{m}\Omega_{s}.$
@@
$g^{(2)}(0)$ $\approx\frac{2\left|C_{g^{\prime},2}\right|^{2}}{\left|C_{g^{\prime},1}\right|^{4}}=\frac{2\left|\eta_{1}\right|^{2}\cdot\left|\chi\right|^{2}}{\left|\eta_{2}\right|^{2}\left|D\right|^{4}}$ (8)
@@
The population in the state $|g^{\prime},1\rangle$ can be optimized by
$\Delta_{s}\Delta_{m}$ $\approx\tilde{\Delta}_{s}\tilde{\Delta}_{m}=\frac{1}{4}g_{ms}^{2}$ (9)
III.2 Mechanisms of The Magnon Blockade
Figure 2: (a) Energy spectrum of the skyrmion qubit-magnon system, showing vacuum Rabi splitting induced by magnon-qubit coupling. (b) Transition pathways in the skyrmion qubit-YIG system.
To explore the mechanism of magnon blockade in a skyrmion-qubit-magnon system, we compute the eigenfunctions and examine the energy spectrum. The ground state is $|g^{\prime},0\rangle$ with eigenvalue $E_{0}=0$, while for $n\geq 1$, the excited (dressed) states are expressed as
$|\psi_{n,\pm}\rangle=C_{g^{\prime},n}|g^{\prime},n\rangle+C_{e^{\prime},n-1}|e^{\prime},n-1\rangle,$
where the normalized coefficients are
$C_{g^{\prime},n}=\pm\frac{g_{ms}\sqrt{n}}{\sqrt{g_{ms}^{2}n+4[n\omega_{m}-E_{n,\pm}]^{2}}},$
and
$C_{e^{\prime},n-1}=\pm\sqrt{\frac{4[n\omega_{m}-E_{n,\pm}]^{2}}{g_{ms}^{2}n+4[n\omega_{m}-E_{n,\pm}]^{2}}}.$
The associated eigenvalues for $\psi_{n,\pm}$ are given by
$E_{n,\pm}=\frac{(2n-1)\omega_{m}+E_{z}}{2}\pm\frac{\sqrt{[\omega_{m}-E_{z}]^{2}+g_{ms}^{2}n}}{2}.$
These coefficients and eigenvalues describe the dressed states in the system and ensure their normalization.
Under the strong-coupling regime, the real parts of the eigenvalues provide an accurate description of the energy spectrum, while the imaginary components, associated with dissipation, can be considered negligible del Valle et al. (2009).
For the case ($\omega_{m}=E_{z}$), the eigenvalues are $E_{n,\pm}=n\omega_{m}\pm\frac{g_{ms}\sqrt{n}}{2}$, and the eigenstates are $|\psi_{n,\pm}\rangle=\frac{1}{\sqrt{2}}|g^{\prime},n\rangle\pm\frac{1}{\sqrt{2}}|e^{\prime},n-1\rangle$. The physical mechanism of the magnon blockade can be explained by the energy-level structure shown in Fig. 2(a). Transitions between the states $|g^{\prime},0\rangle$ and |±$|\pm^{\prime},1\rangle$ occur when the detuning satisfies $\omega_{m}-\omega_{d}=\pm\Delta$ ($\Delta=\frac{1}{2}g_{ms}$). Conversely, transitions from |±$|\pm^{\prime},1\rangle$ to |±$|\pm^{\prime},2\rangle$ are suppressed due to the qubit-induced vacuum Rabi splitting, which requires the condition $2\omega_{m}\pm 2\Delta\neq 2\omega_{m}\pm\frac{\sqrt{2}}{2}g_{ms}$. This suppression arises from the anharmonicity in the energy spectrum caused by the vacuum Rabi splitting. The anharmonicity prevents the second magnon excitation from resonantly reaching the higher states |±$|\pm^{\prime},2\rangle$, confining the system to the single-magnon regime and leading to the CMB effect. In Fig. 2(b), the transition pathways under the combined influence of the probe field and the drive field are depicted. The probe field primarily induces transitions between magnon states, such as $|g^{\prime},0\rangle\xrightarrow{\Omega_{d}}|g^{\prime},1\rangle\xrightarrow{\sqrt{2}\Omega_{d}}|g^{\prime},2\rangle$. Meanwhile, the drive field facilitates transitions between magnon states and qubit states through coupling mechanisms, for example, $|g^{\prime},0\rangle\xrightarrow{\frac{\Omega_{s}}{2}}|e^{\prime},0\rangle\xrightarrow{2^{-1}g_{ms}}|g^{\prime},1\rangle\xrightarrow{\frac{\Omega_{s}}{2}}|e^{\prime},1\rangle\xrightarrow{2^{-1/2}g_{ms}}|g^{\prime},2\rangle$. These pathways may interfere coherently, suppressing the population of the quantum state $|g^{\prime},2\rangle$ and leading to the emergence of the UMB effect.
III.3 Numerical Simulation of the Second-Order Correlation Function
@@
Here, $\kappa_{s}$ and $\kappa_{m}$ denote the decay rates of the skyrmion qubit and magnon mode, with $\kappa_{s}=\kappa_{m}$ assumed for simplicity. The Pauli operators $\tilde{\sigma}_{+}=|\psi_{2}\rangle\langle\psi_{1}|$ and $\tilde{\sigma}_{-}=|\psi_{1}\rangle\langle\psi_{2}|$ are defined using the skyrmion qubit eigenstates $|\psi_{1}\rangle$ and $|\psi_{2}\rangle$. The system density matrix is $\rho$, and the Lindblad superoperators are defined below
@@
$g^{(2)}(0)=\frac{\langle m^{\dagger}m^{\dagger}mm\rangle}{\langle m^{\dagger}m\rangle^{2}}.$ (17)
If $g^{(2)}(0)>1$, this indicates that the steady-state magnon number distribution follows super-Poissonian statistics, exhibiting magnon bunching. Conversely, if $10<g^{(2)}(0)<$ 1, the magnon number distribution satisfies sub-Poissonian statistics, known as magnon antibunching Rabl (2011); Liu et al. (2010); Zheng et al. (2019). Specifically, when $g^{(2)}(0)\rightarrow 0$, magnon blockade occurs.
@@
Figure 3: The steady-state equal-time second-order correlation function $\log_{10}g^{(2)}(0)$ varies with the driving detuning $\Delta/\gamma$ and the coupling strength $g_{ms}/\gamma$ in (a), and with the driving field $\Omega_{s}/\gamma$ and the probe field $\Omega_{d}/\gamma$ in (b). Frequencies are normalized to $\gamma=2\pi\times 1$ MHz for consistency. In (a), the parameters are $\Omega_{s}/\gamma=0.06$ and $\Omega_{d}/\gamma=0.01$, while in (b), $\Delta/\gamma=9.8$ and $g_{ms}/\gamma=19.6$. The other parameters are $\kappa_{m}/\gamma=\kappa_{s}/\gamma=1$ and $n_{th}=0$.
@@
To verify the optimal condition in Eq. (9), we numerically solve the master equation (14) by the Python package QuTiP Johansson et al. (2013, 2012). In Fig. 3(a), the steady-state second-order correlation function $\log_{10}g^{(2)}(0)$ depends on the drive detuning $\Delta/\gamma$ and the coupling strength $g_{ms}/\gamma$ between the magnon and the skyrmion qubit. The transition of $g^{(2)}(0)$ from strong quantum correlations to classical correlations is illustrated by the color gradient, which ranges from deep purple to yellow. The system shows pronounced antibunching near the white dashed lines, where $\Delta/\gamma\approx\pm\frac{1}{2}g_{ms}/\gamma$, and $g^{(2)}(0)$ reaches its minimum. This result highlights the emergence of the MB effect, aligning with the analytically derived the optimal condition. Fig. 3(b) provides a detailed analysis of $g^{(2)}(0)$ as a function of the driving field intensity $\Omega_{s}/\gamma$ and the probe field intensity $\Omega_{d}/\gamma$. As $\Omega_{s}/\gamma$ and $\Omega_{d}/\gamma$ increase, $g^{(2)}(0)$ exhibits significant changes and reaches its minimum near the white dashed lines. The results indicate that the ratio $\Omega_{s}/\Omega_{d}\approx 6$ corresponds to the optimal condition for realizing the MB effect.
@@
Figure 4: The steady-state $\log_{10}g^{(2)}(0)$ is plotted as a function of (a) the coupling strength $g_{ms}$ and (b) the driving field intensity $\Omega_{s}/\gamma$, as well as the dissipation rate $\kappa$, under the condition $\Delta/\gamma\approx\frac{1}{2}g_{ms}/\gamma$. In (a), $\Omega_{s}/\gamma=0.06$ and $\Omega_{d}/\gamma=0.01$, while in (b) $g_{ms}/\gamma=19.6$, and $n_{th}=0$.
Both the probe field and the driving field play a crucial role in enhancing the MB effect. In Fig. 4, we investigate the second-order correlation function $g^{(2)}(0)$ as a function of the coupling strength $g_{ms}$, driving field intensity $\Omega_{s}$, and dissipation rate $\kappa$. In Fig. 4(a), $\log_{10}g^{(2)}(0)$ remains extremely low (deep purple) under conditions of high $g_{ms}$ and low $\kappa$, demonstrating a pronounced antibunching effect. With $g_{ms}/\gamma=19.6$ fixed, Fig. 4(b) explores the influence of $\Omega_{s}$ and $\kappa$ on $g^{(2)}(0)$. For $\Omega_{d}/\gamma=0.01$ and $\Omega_{s}/\gamma\approx 0.06$, $\log_{10}g^{(2)}(0)$ achieves its minimum within the dissipation range $\kappa/\gamma$ between 0.1 and 0.6, corresponding to the strongest MB effect. This confirms that $\Omega_{s}/\Omega_{d}\approx 6$ is optimal for $\kappa/\gamma$ in the range 0.1 to 1.2. These results provide the key parameter values for $\Omega_{s}$, $\kappa$, and $\Omega_{d}$ that maximize the MB effect.
@@
Figure 5: The equal-time second-order correlation function $g^{(2)}(0)$ is shown as a function of the normalized detuning $\Delta/\gamma$. As illustrated in (a), for the magnon drive field intensity $\Omega_{d}/\gamma=0.004$, 0.01, and 0.012, the Skyrmion drive field $\Omega_{s}/\gamma$ is kept constant at 0.06. In (b), for the skyrmion drive field intensity $\Omega_{s}/\gamma=0.001$, 0.05, and 0.09, the magnon drive field $\Omega_{d}/\gamma$ is fixed at 0.01. The coupling strength between the magnon and skyrmion modes is $g_{\text{ms}}/\gamma=19.6$. The other system parameters are $\kappa_{m}/\gamma=\kappa_{q}/\gamma=0.15$, and $n_{th}=0$.
Next, we observe a significant dependence of the second-order correlation function $g^{(2)}(0)$ on the driving detuning $\Delta/\gamma$, as shown in in Fig. 5. In Fig. 5(a), the variation of $g^{(2)}(0)$ is shown for different magnon driving field intensities ($\Omega_{d}/\gamma=0.004,0.01,0.012$), with corresponding ratios $\Omega_{s}/\Omega_{d}=15,6,5$, while the skyrmion driving field intensity is fixed at $\Omega_{s}/\gamma=0.06$. Around $\Delta/\gamma\approx 9.8$, $g^{(2)}(0)$ reaches $10^{-7.3}$, indicating a pronounced antibunching effect and a high-purity magnon blockade state. However, around $\Delta/\gamma\approx-9.8$, the antibunching effect becomes less sensitive to modulation, and $g^{(2)}(0)$ reaches $10^{-3.3}$. In Fig. 5(b), $g^{(2)}(0)$ is plotted as a function of the skyrmion driving field intensity ($\Omega_{s}/\gamma=0.001,0.05,0.09$), with corresponding ratios $\Omega_{s}/\Omega_{d}=0.1,5,9$, while the magnon driving field intensity is fixed at $\Omega_{d}/\gamma=0.01$. In this parameter regime, the MB effect is present but weaker, with $g^{(2)}(0)$ reaching $10^{-2.7}$, $10^{-3.2}$, and $10^{-3.4}$, respectively. Notably, in both Fig. 5(a) and Fig. 5(b), the numerical and analytical results agree almost perfectly, demonstrating the reliability of the model and theoretical calculations. Further analysis confirms that the optimal conditions are $\Omega_{s}/\Omega_{d}\approx 6$, $\Delta/\gamma\approx\pm\frac{1}{2}g_{ms}/\gamma$.